OpenStudy (anonymous):

A flask with 2.50 L of gas is under 3.00 atm of pressure. the gas is heated and the volume increases to 5.00 L. what is work done by the gas? (in J)

OpenStudy (aravindg):

is temp constant??

OpenStudy (anonymous):

i think so , it says nothing about temp

OpenStudy (aravindg):

wel i suppose then it is isothermal expansion

OpenStudy (aravindg):

use the formula for work done in osthermal expansion to get the result

OpenStudy (anonymous):

is it, deltaH=deltaE + deltanRT

OpenStudy (anonymous):

or deltaE = q+w?

OpenStudy (aravindg):

oh not that

OpenStudy (aravindg):

use the formula :below http://en.wikipedia.org/wiki/Isothermal_process

OpenStudy (aravindg):

under calculation of work done

OpenStudy (anonymous):

w=pdv?

OpenStudy (aravindg):

nooo!!

OpenStudy (aravindg):

look under calculation of work done

OpenStudy (anonymous):

i dont have mol

OpenStudy (aravindg):

oh srry then i suppose u jst need to do this: work=PdeltaV

OpenStudy (aravindg):

u get that?

OpenStudy (anonymous):

can u solve it and show me

OpenStudy (aravindg):

srry i am a bit busy

OpenStudy (jfraser):

work done at constant temp is \[w = -P \Delta V\]

OpenStudy (anonymous):

how do i use this equation

OpenStudy (anonymous):

Just plug the numbers in. \[w=-3.00atm(5.00L-2.50L)\]

OpenStudy (anonymous):

the answer is suppose to be 759.8J , but i think ur right and the worksheet may be wrong

OpenStudy (anonymous):

well right now it's in units of liters times atms. We convert that using the equation\[{101.31Joules \over L* atm} \times 7.5 L* atm = 759.8J\] the 101.31 number you should have memorized.

OpenStudy (anonymous):

Always assume a work sheet is right. 99% of the time they are.

OpenStudy (anonymous):

lol thank you so much!

OpenStudy (aravindg):

hmm..... isnt that what i just said :P

OpenStudy (anonymous):

SO SIMPLE QUESTION

OpenStudy (anonymous):

WORK OF EXPANSION=pRESSURE*CHANGE IN VOLUME NOW ACCORDING TO UR QUESTION W=P DELTA V W=3(5-2.50) W=3*2.5 W=7.5